# Network of velocity-coherent filaments formed by supersonic turbulence in a very-high-velocity HI cloud

### v1

@@ 1. FAST observation and data reduction
The observations were conducted in on-the-fly (OTF) mode using the Five-hundred-meter Aperture Spherical radio Telescope (FAST)55. The illuminated aperture () of FAST is 300 meter for zenith angle 26.4∘, and decreases by one third the value for 56. The tracking accuracy is 0.2′. The -band (1–1.5 GHz) front end is a 19-beam dual-polarization (XX and YY) receiver (Supplementary Fig. 1). It has an FWHM beam size () of 3′ at 1.4 GHz, a system temperature of less than 24 K 56, and a beam efficiency for extended sources of 0.85 56, 57.
The spectral backend provides 1024k (about one million) channels, providing a velocity resolution of 0.1 km s-1 at 1.4 GHz, much better than that of the HI4PI survey ( km s-1)36. The analog-to-digital converter (ADC) of the FAST samples the data every 4 ns. The time intervals related to the spectral backend are thus always integer multiples of = s. For our observations, the spectra were recorded every s. For calibration, the system periodically injects noise of K every (with s of noise on and s of noise off).
@@
Split HI cube: First-order polynomial fitting has been adopted to subtract the spectral baseline of the Hi cube. The spectral baseline is mainly composed of the system temperature () and the continuum emission (). We applied 2-d polynomial fitting on the point-source-free regions of the spectral baseline map to further separate and . The resultant is typically 19 K.
Robust check: The continuum image is dominated by background quasars (Supplementary Fig. 2). We applied 2- Gaussian fitting to these point-like sources with peak intensities larger than 3 K, yielding an FWHM of typically 3–3.4 arcmin, comparable to the beam size. The locations and fluxes of the point-like sources obtained by FAST (this work) and the NRAO VLA Sky Survey (NVSS)58 are consistent with each other (Supplementary Fig. 2), with a flux deviation smaller than several percent. We smoothed the Hi cube of FAST (this work) and of HI4PI 36 to the same resolution, and compared the spectra at a randomly chosen location (Supplementary 3). The two spectra are nearly identical. We conclude that the Hi cube we obtained is well calibrated and gridded.
@@ 2. Gaussian decomposition and filament extraction
Gaussian decomposition: We decomposed the velocity subcomponents of the G165 VHVC using scousepy, a widely used multi-component spectral line decomposition tool59, 60. At each pixel, G165 VHVC can be at most divided into three velocity subcomponents (Fig. 1). This procedure decomposed the ppv cube into multiple layers of the fitted (RA, DEC, ) points (Fig. 2).
Algorithm and code of ClimbRidge: ClimbRidge is a Python tool we have written to extract the filaments in -dimensional (-d) space following a gradient-ascent method – the density ridge algorithm50, 51. This algorithm models filaments in n- space as density ridges, where the eigenvector of the maximum eigenvalue of the Hessian matrix is parallel to the direction of the gradient. In practice, to find out the density ridges, a set of points are scattered in the -d space, and each point climbs along the gradient in the subspace of the other n-1 eigenvectors (those of all the eigenvalues except the maximum one). We provide both the CPU and GPU versions of ClimbRidge.
Test of ClimbRidge in 2-d case: In 2-d cases, the data at each pixel of the integrated intensity map are taken as the input of ClimbRidge to construct the density ridges. For comparison, we also apply FilFinder61 to find filaments on the 2-d map (Extended Fig. 3). The filaments found by the two methods are similar, and ClimbRidge is more sensitive to weak structures. FilFinder is a 2-d tool, while ClimbRidge can be directly applied to data of any dimension.
@@
Extract filaments in 3-d ppv cube: In 3-d cases, the fitting results of scousepy are taken as the input of ClimbRidge. We reconstructed a smoothed ppv cube through convolving the fitting results of scousepy with a Gaussian kernel. The FWHM of the Gaussian kernal is 3.5′, 3.5′ and 15 km s-1 in the RA, DEC and velocity direction, respectively. In total, 15 000 points are scattered in the ppv space and move along the projected gradient. The resultant density ridges are shown in Fig. 1 and Extended Data Fig. 4.
@@
MST analysis: We have developed a tool, fil3d, to analyse the ridge points produced by ClimbRidge, including (1) trimming out the isolated points, (2) minimal spanning tree analysis, (3) and find out the shortest path between two given points. We have built the minimal spanning trees (MSTs) to connect the density-ridge points (Extended Data Fig. 6), using the public Python package MiSTree, for both the 2-d and 3-d cases. To build the 3-d tree, the Euclidean distance is normalized by the units of (pixel, pixel, km s-1) in the three directions. A segment is defined as the path that connects two neighbouring nodes. A branch is defined as one or more connected segments. We also redistributed the density-ridge points through randomly rearranging their RA, DEC and , and then built a random tree. The maximuum segment length of the random tree is obviously shorter than that of the 3-d tree (Fig. 3). The 2-d and 3-d trees of G165 VHVC are qualitatively similar to each other (Extended Data Fig. 6). The lower end of the segment-length PDF of the 3-d tree is more flattened than that of the 2-d tree. It implies that the density ridges in the ppv cube can trace more short filamentary branches, e.g., in the crowded region of branch B3 (Extended Data Fig. 4).
3. Derive physical parameters Adopting a thickness of the G165 VHVC of 1.5∘ (the value comparable with its angular size) and a layer number of 4, the thickness of one layer () is estimated to be
@@
The Hi emission of the G165 VHVC is weak and optically thin (Fig. 1). The Hi column density can be directly calculated through 62
@@
The median value of the Hi integrated intensity is 37 K km s-1, yielding a median column density of cm-2. It corresponds to a typical global volume density, , of 0.2 cm-2. The linear mass () of the Hi filament can then be derived through
where is the distance between G165 VHVC and us. The gravity-related critical density of a filament can be estimated through 7
@@
where is the gravitational constant. Inserting the typical values of Table 1 into Eq. 5, we get pc-1. Taking as 10 km s-1, we get pc-1, much larger than the value of . The total Hi flux of the G165 VHVC is 350 K km s -1 deg2, corresponding to a total mass of
@@
adopting and kpc. is also much larger than , implying that the self-gravity of G165 VHVC can be ignored.
@@ 4. Ideal MHD simulation
We use the multi-physics code ORION2 that is based on the Godunov scheme 63, 64 to simulate the observed G165 VHVC. The ideal MHD module in ORION2 is based on the dimensionally unsplit corner transport upwind scheme65 incorporating the constrained transport framework66, 67. ORION2 also has the adaptive refinement capability but we shall only use a single grid in our simulations here. From observation, the G165 VHVC is likely supersonic with a of 3. Besides this value, we have also performed a few more models at of 0.5 (subsonic), 1.0 (transonic), and 4.0 (even more supersonic) for comparison. We do not have magnetic field information of the region but we expect a magnetic field will affect the ionized WNM. Therefore, we perform a parameter study of with values 0.5, 1, and 2 using turbulence boxes with periodic boundary conditions. Here, with the Alfvén speed km s-1 for an Hi system. We assume that the VHVC is exposed to a uniform ultraviolet field and has a constant temperature. The overall volume density of G165 VHVC is smaller than the typical phase-transition density of WNM 43. The enhanced cooling rate in the most central regions of the filamentary hubs may lead to the existence of a colder HI phase, but it should not significantly influence the global dynamics. Therefore, we use the isothermal equation of state in the simulation. As discussed above, we also ignore gravity in the simulations. The simulated models are thus scaleless and only controlled by and . In each model, we start with an initially uniform density and magnetic field along the z-axis. We use the recipe of MacLow68 in driving the turbulence boxes with a pure solenoidal velocity field continuously at the largest scale (k=1-2) for two crossing times , where is the length of the box and is the root-mean-square turbulence velocity. For , is the sound speed . For comparison with observation, we only use the simulation data at .
@@
Data Availability Owing to data size, the reduced data used for this study are available from the corresponding authors upon reasonable request.
@@
Acknowledgements We wish to thank the staff of the FAST for their kind help provided during the observations. This work has been supported by the National Key R&D Program of China (No. 2022YFA1603100) and National Natural Science Foundation of China (NSFC) through grants (No. 12203086, No.12073061, No.12122307 and No. 12033005). X.L. acknowledges the supports by CPSF No. 2022M723278. T.L. and P.S.Li acknowledge the supports by the Tianchi Talent Program of Xinjiang Uygur Autonomous Region and the international partnership program of Chinese Academy of Sciences through grant No.114231KYSB20200009. This research was carried out in part at the Jet Propulsion Laboratory, which is operated by the California Institute of Technology under a contract with the National Aeronautics and Space Administration (80NM0018D0004). We thank the anonymous reviewers for providing feedback that helps to solidify this work.
@@
Filament | length | V | |||
(arcmin) | (arcmin) | (km s-1) | (km s-1 pc-1) | (K) | |
B1 | 90 | 6.1 | 25.3 | -0.02 | 0.73 |
B2 | 54 | 9.5 | 23.9 | 0.1 | 0.85 |
B3 | 60 | 7.0 | 29.4 | -0.1 | 0.40 |
Note: The listed parameters include the angular length, velocity gradient (), angular width (), velocity width (), and the averaged brightness temperature (). Adopting a distance of 15 kpc, the converting factor between the linear scale (pc) and angular scale (arcmin) is 4.36 pc arcmin-1.

### v2

@@ 1. FAST observation and data reduction
The observations were conducted in on-the-fly (OTF) mode using the Five-hundred-meter Aperture Spherical radio Telescope (FAST)61. The illuminated aperture () of FAST is 300 meter for zenith angle 26.4∘, and decreases by one third the value for 62. The tracking accuracy is 0.2′. The -band (1–1.5 GHz) front end is a 19-beam dual-polarization (XX and YY) receiver (Supplementary Fig. 2). It has an FWHM beam size () of 3′ at 1.4 GHz, a system temperature of less than 24 K 62, and a beam efficiency for extended sources of 0.85 62, 63.
The spectral backend provides 1024k (about one million) channels, providing a velocity resolution of 0.1 km s-1 at 1.4 GHz, much better than that of the HI4PI survey ( km s-1)41. The analog-to-digital converter (ADC) of the FAST samples the data every 4 ns. The time intervals related to the spectral backend are thus always integer multiples of = s. For our observations, the spectra were recorded every s. For calibration, the system periodically injects noise of K every (with s of noise on and s of noise off).
@@
Split HI cube: First-order polynomial fitting has been adopted to subtract the spectral baseline of the Hi cube. The spectral baseline is mainly composed of the system temperature () and the continuum emission (). We applied 2D polynomial fitting on the point-source-free regions of the spectral baseline map to further separate and . The resultant is typically 19 K. The Hi cube of G165 VHVC can be found in an online repository 64.
Robust check: The continuum image is dominated by background quasars (Supplementary Fig. 3). We applied two-dimensional Gaussian fitting to point-like sources with peak intensities greater than 3 K, yielding full width at half maximum (FWHM) values typically in the range of 3–3.4 arcmin, comparable to the beam size. The point-like sources identified by FAST, which has a larger beam size, encompass those detected by the NRAO VLA Sky Survey (NVSS) 65. The continuum fluxes of the strong point-like sources measured by FAST and NVSS are consistent (Supplementary Fig. 3), with flux deviations typically smaller than six percent. We smoothed the Hi cube from FAST (this work) and from HI4PI 41 to the same resolution, and compared the spectra at a randomly selected location (Supplementary Fig. 4). The two spectra are nearly identical. We conclude that the Hi cube we obtained is well calibrated and properly gridded.
@@ 2. Gaussian decomposition and filament extraction
Gaussian decomposition: We decomposed the velocity subcomponents of the G165 VHVC using scousepy, a widely used multi-component spectral line decomposition tool66, 67. The velocity range for fitting is from to km s-1. At each pixel, G165 VHVC can be at most divided into three velocity subcomponents (Fig. 1). This procedure decomposed the ppv cube into multiple layers of the fitted (RA, DEC, ) points (Fig. 2). The residuals of the multiple Gaussian fitting are shown in Supplementary Fig. 5.
Algorithm and code of ClimbRidge: ClimbRidge is a Python tool we have written to extract the filaments in -dimensional (-d) space following a gradient-ascent method – the density ridge algorithm55, 56. This algorithm constructs a smooth surface (or “potential”) by convolving a set of input data points with Gaussian kernels:
Here, is the weight of the input point, is its location along the dimension, and is the dispersion (corresponding to FWHM/) of the Gaussian kernel along the axis. The algorithm then identifies the density ridges of the surface (), defined as the locations where the eigenvector corresponding to the largest eigenvalue of the Hessian matrix of is aligned with the gradient of . These density ridges are interpreted as filaments traced by the input data points. To numerically locate the density ridges, a set of test points is scattered throughout the -dimensional space. Each test point ascends the surface by following the gradient direction projected onto the subspace spanned by the eigenvectors corresponding to the smaller eigenvalues of the Hessian matrix of . We provide both CPU and GPU implementations of ClimbRidge.
Test of ClimbRidge in 2D case: In the 2D case, the Gaussian fitting results (RA, DEC, and integrated intensity) from scousepy are used as the input points for ClimbRidge to construct the 2D density ridges. To generate a smooth surface from the input data, a Gaussian kernel with a FWHM of 3.5′ in both the RA and DEC directions is applied, and the fitted integrated intensity is adopted as the weight of each input point. The FWHM of the Gaussian kernel is chosen to be comparable to the beam size of FAST, in order to suppress noise that could cause the test points to become trapped in spurious local maxima, while still preserving sufficient spatial resolution. For comparison, we also use FilFinder 68 to detect filaments on the integrated intensity map (Extended Fig. 3). The filaments identified by both methods are similar, with ClimbRidge exhibiting greater sensitivity to weak structures. While FilFinder is limited to 2D analysis, ClimbRidge can be directly applied to data of arbitrary dimensionality.
@@
Extract filaments in 3D ppv cube: In the 3D case, the Gaussian fitting results (RA, DEC, velocity, and integrated intensity) from scousepy are used as the input points for ClimbRidge. The FWHM of the Gaussian kernel is adopted to be 3.5′ in both the RA and DEC directions, and 15 km s-1 in the velocity direction. Here, a FWHM of 15 km s-1, slightly smaller than the typical line width, is adopted to ensure that the scattered points climb toward different velocity layers. A total of 15,000 test points are scattered throughout the ppv space and ascend along the projected gradient. The resulting density ridges are shown in Fig. 1 and Extended Data Fig. 4.
@@
MST analysis: We have developed a tool, fil3d, to analyse the ridge points produced by ClimbRidge, including (1) trimming out the isolated points, (2) minimal spanning tree analysis, (3) and find out the shortest path between two given points. We have built the minimal spanning trees (MSTs) to connect the density-ridge points (Extended Data Fig. 5), using the public Python package MiSTree, for both the 2D and 3D cases. To build the 3D tree, the Euclidean distance is normalized by the units of (pixel, pixel, km s-1) in the three directions. A segment is defined as the path that connects two neighbouring nodes. A branch is defined as one or more connected segments. We also redistributed the density-ridge points through randomly rearranging their RA, DEC and , and then built a random tree. The maximuum segment length of the random tree is obviously shorter than that of the 3D tree (Fig. 3). The 2D and 3D trees of G165 VHVC are qualitatively similar to each other (Extended Data Fig. 5). The lower end of the segment-length PDF of the 3D tree is more flattened than that of the 2D tree. It implies that the density ridges in the ppv cube can trace shorter filamentary branches, especially in crowded regions such as branch B3 (Extended Data Fig. 4). This may be because, in the 3D case, dynamically unrelated branches that appear connected in projection can be more effectively disentangled.
3. Derive physical parameters
There are three main velocity layers, with flare gas extending to velocities greater than 240 km s-1 (Extended Data Fig. 4). Five velocity groups are fitted from the central velocity distribution (Fig. 3). Flare gas contributes to a number of velocity groups in addition to the main velocity layers. Therefore, we adopt an effective layer number of 4 (the mean of the numbers of main velocity layers and velocity groups) to estimate the layer thickness. Adopting a total thickness of the G165 VHVC of 1.5∘ (a value comparable to its angular size), the layer thickness () is roughly estimated to be
@@ 3. Derive physical parameters
where is the distance between G165 VHVC and us. The Hi emission of the G165 VHVC is weak and optically thin (Fig. 1). The Hi column density can be directly calculated through 69, 2
@@
where is the FWHM velocity width and is the averaged peak brightness temperature. Note that we have accounted for the conversion factor of 1.064 from to the integrated intensity of a Gaussian profile. The median value of the Hi integrated intensity () is 37 K km s-1, yielding a median column density of cm-2. It corresponds to a typical global volume density, , of 0.2 cm-2. The linear mass () of the Hi filament can then be derived through
where is the FWHM angular width of the filament.
The gravity-related critical density of a filament can be estimated through 7
@@
where is the gravitational constant, and is the velocity dispersion (FWHM/). Inserting the typical values of Extended Data Table 1 into Eq. 6, we get pc-1. Taking as 10 km s-1, we get pc-1, much larger than the value of . The total Hi flux of the G165 VHVC is 350 K km s -1 deg2, corresponding to a total mass of
@@
Here, is the physical size of G165 VHVC, estimated to be . is also much larger than , implying that the self-gravity of G165 VHVC can be ignored.
@@ 4. Ideal MHD simulation
We use the multi-physics code ORION2 that is based on the Godunov scheme 70, 71 to simulate the observed G165 VHVC. The ideal MHD module in ORION2 is based on the dimensionally unsplit corner transport upwind scheme72 incorporating the constrained transport framework73, 74. ORION2 also has the adaptive refinement capability but we shall only use a single grid in our simulations here. From observation, the G165 VHVC is likely supersonic with a of 3. Besides this value, we have also performed a few more models at of 0.5 (subsonic), 1.0 (transonic), and 4.0 (even more supersonic) for comparison. We do not have magnetic field information of the region but we expect a magnetic field will affect the ionized WNM. Therefore, we perform a parameter study of with values 0.5, 1, and 2 using turbulence boxes with periodic boundary conditions. Here, with the Alfvén speed km s-1 for an Hi system. We assume that the VHVC is exposed to a uniform ultraviolet field and has a constant temperature. The overall volume density of G165 VHVC is smaller than the typical phase-transition density of WNM 48. The enhanced cooling rate in the most central regions of the filamentary hubs may lead to the existence of a colder HI phase, but it should not significantly influence the global dynamics. Therefore, we use the isothermal equation of state in the simulation. As discussed above, we also ignore gravity in the simulations. The simulated models are thus scaleless and only controlled by and . In each model, we start with an initially uniform density and magnetic field along the z-axis. We use the recipe of MacLow75 in driving the turbulence boxes with a pure solenoidal velocity field continuously at the largest scale (k=1-2) for two crossing times , where is the length of the box and is the root-mean-square turbulence velocity. For , is the sound speed . For comparison with observation, we only use the simulation data at .
@@
Data Availability The FAST Hi cube of G165 VHVC, smoothed to a velocity resolution of 1 km s-1, can be found at https://zenodo.org/records/15347725. Owing to data size, the full data used for this study are available from the corresponding authors upon reasonable request.
@@
Acknowledgements We wish to thank the staff of the FAST for their kind help provided during the observations. X.L. and T.L. have been supported by the National Key R&D Program of China (No. 2022YFA1603100). X.L. has also been supported by the Strategic Priority Research Program of the Chinese Academy of Sciences under Grant No. XDB0800303. T.L. acknowledges National Natural Science Foundation of China (NSFC) No. 12017061 and 12122307. T.L. also acknowledges the supports by the Tianchi Talent Program of Xinjiang Uygur Autonomous Region and the PIFI program of Chinese Academy of Sciences through grant No. 2025PG0009. P.L. acknowledges NSFC No. 1241101426. S.Q. acknowledges NSFC No.12033005. N.T. acknowledges the supports by the NSFC No. 12473023, and by the University Annual Scientific Research Plan of Anhui Province (No. 2023AH030052 and 2022AH010013). This research was carried out in part at the Jet Propulsion Laboratory, which is operated by the California Institute of Technology under a contract with the National Aeronautics and Space Administration (80NM0018D0004). We thank the anonymous reviewers for providing feedback that helps to solidify this work.
@@
Extended Data figures
Extended Data tables
![[Uncaptioned image]](extracted/6516493/ED_Table1.jpg)
Note: The listed parameters include the angular length, velocity gradient (), FWHM angular width (), FWHM velocity width (), and the averaged peak brightness temperature () of the typical filaments (Extended Data Fig. 6). Adopting a distance of 15 kpc, the converting factor between the linear scale (pc) and angular scale (arcmin) is 4.36 pc arcmin-1.
Supplementary figures